Solve the equation using factoring: 49p^2 – 112p + 64 = 0
Answer by Alan3354(8229)

Solve the equation using factoring: 49p^2 – 112p + 64 = 0
(7p-8)^2 = 0
7p-8 = 0
7p = 8
p = 8/7